a triangle has sides that measures 15 cm, 20 cm, and 25 cm. find the exact volume of the solid of revolution formed when the triangle is revolved about the side of length 25 cm.


1
Expert's answer
2021-11-23T09:42:48-0500

The triangle is right angled

Solid of revolution will be a bicone



Using Hero's formula


Area if triangle A"=\sqrt{S(s-a)(s-b)(s-c)}"


Where a,b,c are sides of the triangle and S "=\frac{1}{2}(a+b+c)"



"A=\sqrt{30(30-15)(30-25)(30-20)}"


"=150cm^2"



Height, h of the triangle with "25cm" as base is given by;


"\frac{1}{2}×25×h=150"


"h=12cm"



The solid of revolution formed is a bicone







Volume of cone "=\frac{1}{3}\pi\>r^2h" where


"r=" Radius

h= height of cone


Volume of the bicone ;


"\frac{1}{3}\pi×12^2(h_1)+\frac{1}{3}\pi×12^2(h_2)"



"=\frac{1}{3}\pi×12^2(h_1+h_2)"



"=\frac{1}{3}\pi×12^2×25"


"=3769.9cm^2"